OpenStudy (anonymous):

Solve: x2 - 20x + 36 = 0

OpenStudy (callisto):

x^2 - 20x +36 = 0 x^2 -2x - 18x +36 = 0 x(x-2) -18(x-2) = 0 (x-2)(x-18) = 0 x-2 = 0 or x-18 = 0 <-solve these equations

OpenStudy (mathlegend):

Shouldn't you have walked her through from the beginning?

OpenStudy (callisto):

To solve quadratic equations, you can do it like the above, that is factoring the expression on the left, put each factor = 0 and solve them OR Use quadratic formula.

OpenStudy (mathlegend):

You basically just gave out the answer.

OpenStudy (callisto):

@MathLegend Did I? Where's the answer?

OpenStudy (mathlegend):

I said you basically did if you can read.

OpenStudy (callisto):

The only mistake that I made here was not explain with words But basically, I did show the asker the way to work the answer out.

OpenStudy (mathlegend):

Yes, but my point is.. since she posted this. Don't you think she may need some explanation along the way?

OpenStudy (mathlegend):

Liek as you go from Step 1 to Step 2

OpenStudy (mathlegend):

Like*

OpenStudy (callisto):

If that's what you want, x^2 - 20x +36 = 0 Split the middle term so that you can factor the expression easier x^2 -2x - 18x +36 = 0 Take out the common factor for the first two terms and last two terms x(x-2) -18(x-2) = 0 Take out the common factor again (x-2)(x-18) = 0 Put each factor = 0 x-2 = 0 or x-18 = 0 solve these equations on your own

OpenStudy (mathlegend):

Okay, I am just saying because a Mod told me that when I first joined.

OpenStudy (callisto):

I'm sorry but I still don't think I'm giving the answer.

OpenStudy (mathlegend):

Okay. :)

OpenStudy (anonymous):

@Ghatiii, did you get your answers?

OpenStudy (anonymous):

Or.....Do u knw Quadratic Formuld..)) it will be very easy

OpenStudy (callisto):

Just want to make few points clear. 1. If I was giving answer in the first post, I'm still giving answer in the second time I did it, even with explanations. 2. I didn't intend to give answers, as always. 3. I understand that I should have explained using words, but without using a word doesn't mean that I give answer. At least, I leave some steps for the asker to do it. 4. If you have noticed since I've started typing in this post. After I left the first comment, I actually kept typing again as I found it is better to leave some words to the asker that I'm doing factorization, as you've seen above. I apologize that I didn't explain what I had done in the first post, but I still insist that I'm not giving answers.

OpenStudy (anonymous):